OpenStudy (anonymous):

Find the distance between the points to the nearest tenth. L(10,14), M(-8,14) Find the distance between the points to the nearest tenth. L(10,14), M(-8,14) @Mathematics

OpenStudy (anonymous):

do u know the distance formula? I am trying to figure out how to type it.

OpenStudy (anonymous):

Yeah d=\[\sqrt{(X2-X1)^2 +(Y2-Y1)^2}\]

OpenStudy (anonymous):

I think XD

OpenStudy (anonymous):

so if u set x2 to 10 x1 to 8.. and y2 and y1 to 14 and plug them in

OpenStudy (anonymous):

\[\sqrt{(10-8)^2+(14-14)^2}\]

OpenStudy (anonymous):

Horizontal line - LM=2

OpenStudy (anonymous):

324

OpenStudy (anonymous):

I got 2 too :] Thank you!

OpenStudy (anonymous):

bleh missed the negative on the 8.. sec computing

OpenStudy (anonymous):

18

OpenStudy (anonymous):

?

OpenStudy (anonymous):

u said the second point was (-8, 14) so \[\sqrt{(10-(-8))^2+(14-14)^2}\] thus 10+8 = 18^2 = \[\sqrt{324+0}\] = 18

OpenStudy (anonymous):

Ohh ok thanks!